Conside the linear equation D=RT, where D is distance in miles, R is speed in miles per hours. If I traveled at a speed of 50 miles per hour, find the time it takes me to catch up to a car traveling at 30 miles per hour that had a 200 mile head start.
Answer by scott8148(6628)

when you catch the first car, you will have traveled the same distance ... the time it takes to catch up is your travel time ... this is the time that the first car travels AFTER the 200 mile head start

50*t=30*t+200 ... 20t=200 ... t=10 hours